# On generating mapping class groups by pseudo-Anosov elements

### v2

@@ 2. Preliminaries
(28) $t_{a_{1}}t_{c_{1}}^{-1}=t_{a_{1}}t_{a_{3}}^{-1}\cdot t_{a_{3}}t_{c_{1}}^{-1}\in G_{2},$
(29) $t_{a_{2}}t_{c_{2}}^{-1}=(t_{a_{1}}t_{a_{2}}^{-1})^{-1}\cdot t_{a_{1}}t_{c_{2}}^{-1}\in G_{2}.$
Since it is immediate that $\sigma_{n}(a_{1},b_{1})=(a_{2},t_{c_{2}}(b_{2}))$, by (25) and $\sigma_{n}\in G_{2}$ we have
$G_{2}\ni t_{a_{2}}^{-1}t_{t_{c_{2}}(b_{2})}=t_{a_{2}}^{-1}t_{c_{2}}t_{b_{2}}t_{c_{2}}^{-1}=(t_{c_{2}}^{-1}t_{a_{2}})^{-1}t_{b_{2}}t_{c_{2}}^{-1}.$
By (29) and the disjointness of $a_{2}$ and $c_{2}$, we see that $t_{b_{2}}t_{c_{2}}^{-1}\in G_{2}$. Therefore, by (27), the following holds:
$t_{a_{1}}t_{b_{2}}^{-1}=t_{a_{1}}t_{c_{2}}^{-1}\cdot(t_{b_{2}}t_{c_{2}}^{-1})^{-1}\in G_{2}.$
From (22)–(24), (27)–(29) and this, we have $t_{a_{3}}\in G_{2}$ using Lemma 4.5. By a similar argument to that in the case of $g\geq 4$, we see that $G_{2}=\mathcal{M}_{3}$. ∎
@@
Suppose that $g\geq 4$. By a similar argument to the proof of Theorem 4.2, we see that $t_{a_{1}}t_{a_{2}}^{-1},t_{a_{1}}t_{a_{3}}^{-1},t_{a_{1}}t_{b_{3}}^{-1}\in G_{4}$. Moreover, the element $\rho^{\prime}=rt_{c_{1}}$ is equal to $\sigma_{1}$ in the proof of Theorem 4.2. Therefore, we have $t_{a_{1}}t_{b_{2}}^{-1},$ $t_{a_{1}}t_{c_{1}}^{-1},$ $t_{a_{2}}t_{c_{2}}^{-1},$ $t_{a_{1}}t_{c_{2}}^{-1}\in G_{4}$, and hence the elements $t_{a_{1}},t_{b_{1}},t_{c_{1}},r$ are contained in $G_{4}$. By a similar argument to the proof of Theorem 4.1, we see that $G_{4}=\mathcal{M}_{g}$ for $g\geq 4$. The same reasoning applies to the case $g=3$. This finishes the proof. ∎

### v3

@@ 2. Preliminaries
(30) $t_{a_{1}}t_{c_{1}}^{-1}=t_{a_{1}}t_{a_{3}}^{-1}\cdot t_{a_{3}}t_{c_{1}}^{-1}\in G_{2},\ \ \mathrm{and}$
(31) $t_{c_{2}}^{-1}t_{a_{2}}=t_{a_{2}}t_{c_{2}}^{-1}=(t_{a_{1}}t_{a_{2}}^{-1})^{-1}\cdot t_{a_{1}}t_{c_{2}}^{-1}\in G_{2}.$
By $\sigma_{n}(a_{1},b_{1})=(a_{2},t_{c_{2}}(b_{2}))$ and (27), the element $t_{a_{2}}^{-1}t_{t_{c_{2}}(b_{2})}=t_{a_{2}}^{-1}t_{c_{2}}t_{b_{2}}t_{c_{2}}^{-1}$ is in $G_{2}$. Therefore, the element $t_{b_{2}}t_{c_{2}}^{-1}$ is in $G_{2}$ by (31), and the element $t_{a_{1}}t_{b_{2}}^{-1}=t_{a_{1}}t_{c_{2}}^{-1}\cdot(t_{b_{2}}t_{c_{2}}^{-1})^{-1}$ is in $G_{2}$ by (29). From (23)–(25), (29)–(31) and this, the Dehn twist $t_{a_{3}}$ is in $G_{2}$ using Lemma 4.5. By a similar argument to that in the case of $g\geq 4$, we see that $G_{2}=\mathcal{M}_{3}$. ∎
@@
Suppose that $g\geq 4$. By a similar argument to the proof of Theorem 4.2, we see that $t_{a_{1}}t_{a_{2}}^{-1}$, $t_{a_{1}}t_{a_{3}}^{-1}$ and $t_{a_{1}}t_{b_{3}}^{-1}$ are in $G_{4}$. Moreover, the element $\rho^{\prime}=rt_{c_{1}}$ is equal to $\sigma_{1}$ in the proof of Theorem 4.2. Therefore, $t_{a_{1}}t_{b_{2}}^{-1}$, $t_{a_{1}}t_{c_{1}}^{-1}$, $t_{a_{2}}t_{c_{2}}^{-1}$ and $t_{a_{1}}t_{c_{2}}^{-1}$ are in $G_{4}$, and hence the elements $t_{a_{1}}$, $t_{b_{1}}$, $t_{c_{1}}$ and $r$ are in $G_{4}$. By a similar argument to the proof of Theorem 4.1, we see that $G_{4}=\mathcal{M}_{g}$ for $g\geq 4$. The same reasoning applies to the case $g=3$. This finishes the proof. ∎